Two positive numbers have a difference of 8. The larger number is one more than twice the smaller. Find the two numbers.

Answer by CubeyThePenguin(3113)

x = larger number
y = smaller number

x - y = 8
x = 2y + 1

Use substitution.

2y + 1 - y = 8
y + 1 = 8
y = 7

The two numbers are 15 and 7.